# How Far Is Video Generation from World Model: A Physical Law Perspective

### v1

@@ A.1 More related works
Physical reasoning. It refers to the ability to understand and predict the way objects will interact under certain conditions according to physical laws. Melnik et al. (2023) categorize physical reasoning tasks into passive and interactive tasks. Passive tasks often require the AI to predict certain properties of objects, e.g., materials (Bell et al., 2015; Bouman et al., 2013), physical parameters (Wu et al., 2016), stability of a physical system (Groth et al., 2018), or involve question-answering for the agent to recognize conceptual differences (Weitnauer et al., 2023; Yi et al., 2019), or describe a physical scenario (Ates et al., 2020). For interactive tasks, AI is required to control some objects in the environment to complete certain tasks based on physical commonsense, e.g., solving classical mechanics puzzles (Bakhtin et al., 2019), flying a bird to reach a target position (Xue et al., 2023), and using a tool (Allen et al., 2020). Two works closely related to ours are Girdhar et al. (2020) and de Silva et al. (2020). Girdhar et al. (2020) introduce a forward prediction model to aid physical reasoning but do not address what is learned in the prediction model. de Silva et al. (2020) attempt to discover universal physical laws from data with abstracted internal states and human expertise introduced in the dynamic model design. In contrast, we focus on recovering physical laws from raw observation without any human priors, akin to a newborn baby.
A.2 Diffusion preliminaries
@@
where the target can be the corrupted noise , the original video , or the velocity between data and noise. Following Salimans & Ho (2022), we use the velocity prediction to train the video diffusion models, denoted by
Also, to eliminate the training and inference gap and ensure a zero signal-to-noise ratio at the final timestep, following Lin et al. (2024), we set to when .
A.3 VAE
A.3.1 VAE Architecture and Pretrain
We commence with the structure of the SD1.5-VAE, retaining the majority of the original 2D convolution, group normalization, and attention mechanisms on the spatial dimensions. To inflate this structure into a spatial-temporal auto-encoder, we convert the final few 2D downsample blocks of the encoder and the initial few 2D upsample blocks of the decoder into 3D ones, and employ multiple extra 1D layers to enhance temporal modeling. For all the downsample blocks where temporal downsampling is required, we replace all the original 2D downsample layers with re-initialized causal 3D downsample layers by adding causal paddings to the head of the frame sequence (Yu et al., 2023b) and introduce an additional causal 1D convolution layer after the original ResNetBlock. As for the decoder part, all the 2D Nearest-Interpolation operations are substituted by a 2D convolution layer and a channel-to-space transformation, which are specifically initialized to behave precisely the same as Nearest-Interpolation operations before the first training step. For the discriminator part, we inherit the structure of the original 2D PatchGAN (Isola et al., 2016) discriminator used by SD1.5-VAE, and design a 3D PatchGAN discriminator based on the 2D version. Different from the generator module, we train the group of discriminators from scratch for the consideration of stability. Subsequently, all parameters of the (2+1)D-VAE are jointly trained with high-quality image and video data to preserve the capability of appearance modeling and to enable motion modeling. For the image dataset, we filter data samples from LAION-Aesthetics (Beaumont & Schuhmann, 2022), COYO (Byeon et al., 2022) and DataComp (Gadre et al., 2023) with high aesthetics and clarity to form a high-quality subset. As for the video dataset, we collect a high-quality subset from Vimeo-90K (Xue et al., 2017), Panda-70M (Wang et al., 2020) and HDVG Wang et al. (2023). In the training process, we train the entire structure for 1M steps and only the random resized crop and random horizontal flip are applied in the data augmentation process.
A.3.2 VAE Reconstruction
@@
A.4 Fundamental Physical Scenarios Data
@@
For all scenarios, we filter out videos where the ball exits the field of view prematurely by inspecting the simulation state, as these videos do not provide sufficient meaningful information for the model.
A.5 More Experiments and Discussions
A.5.1 Can Language and Numerics Aid in Learning Physical Laws?
@@
As discussed in Section 3, video generation based solely on image frames fails to learn physical laws, showing significant prediction errors in OOD scenarios, despite containing all the necessary information. In reinforcement learning, numerical values (e.g., states) are often used as conditions for world models (Hafner et al., 2023), and language representations have shown generalization capabilities in LLMs (Riveland & Pouget, 2024). This raises the question: can additional multimodal inputs, such as numerics and text, improve video prediction and capture physical laws? We experimented with collision scenarios and DiT-B models, adding two variants: one conditioned on vision and numerics, and the other on vision and text. For numeric conditioning, we map the state vectors to embeddings and add the layer-wise features to video tokens. For text, we converted initial physical states into natural language descriptions, obtained text embeddings using a T5 encoder, and then add a cross-attention layer to aggregate textual representations for video tokens.
@@
A.5.2 Principle Behind Data Retrieval in the Diffusion Model
@@
Since the diffusion model is trained by minimizing the loss associated with predicting VAE latent, we hypothesize that the prioritization may be related to the distance in VAE latent space (though we use pixel space here for clearer illustration) between the test conditions and the training set. Intuitively, when comparing color and shape as in Figure 9 (1), a shape change from a ball to a rectangle results in minor pixel variation, primarily at the corners. In contrast, a color change from blue to red causes a more significant pixel difference. Thus, the model tends to preserve color while allowing shape to vary. From the perspective of pixel variation, the prioritization of color > size > velocity > shape can be explained by the extent of pixel change associated with each attribute. Changes in color typically result in large pixel variations because it affects nearly every pixel across its surface. In contrast, changes in size modify the number of pixels but do not drastically alter the individual pixels’ values. Velocity affects pixel positions over time, leading to moderate variation as the object shifts, while shape changes often involve only localized pixel adjustments, such as at edges or corners. Therefore, the model prioritizes color because it causes the most significant pixel changes, while shape changes are less impactful in terms of pixel variation.
@@
To further validate this hypothesis, we designed a variant experiment comparing color and shape, as shown in Figure 13. In this case, we use a blue ball and a red ring. For the ring to transform into the ball without changing color, it would need to remove the ring’s external color, turning it into blank space, and then fill the internal blank space with the ball’s color, resulting in significant pixel variation. Interestingly, in this scenario, unlike the previous experiments shown in Figure 9 (1), the prioritization of color > shape does not hold. The red ring can transform into either a red ball or a blue ring, as demonstrated by the examples. This observation suggests that the model’s prioritization may indeed depend on the complexity of the pixel transformations required for each attribute change. Future work could explore more precise measurements of these variations in pixel or VAE latent space to better understand the model’s training data retrieval process.
A.5.3 Failure Cases in Combinatorial Generalization
@@
In Figure 11, we present successful cases of spatial and temporal combinatorial generalization, where the trained model can combine these two types of events across spatial and temporal dimensions to generate videos under unseen conditions. However, the model does not always succeed in performing such compositions, and here we illustrate some failure cases. As shown in Figure 14, when the training set lacks a red ball in a bounce event, the model sometimes struggles. In the model’s generated video, the red ball sometimes disappears after a collision. This likely stems from the model’s data retrieval mechanism: since the training set does not include a red ball in a collision scenario, when collision happens, the model retrieves similar training cases without the red ball, causing it to vanish post-collision. In summary, while combinatorial generalization allows the diffusion model to generate novel videos by composing spatial and temporal segments from the training set, its reliance on data retrieval limits its effectiveness. As a result, the model may produce unrealistic outcomes by retrieving and combining segments without understanding the underlying rules.
A.6 visualization

### v2

@@ Appendix A More related works
Physical reasoning. It refers to the ability to understand and predict the way objects will interact under certain conditions according to physical laws. (Melnik et al., 2023) categorize physical reasoning tasks into passive and interactive tasks. Passive tasks often require the AI to predict certain properties of objects, e.g., materials (Bell et al., 2015; Bouman et al., 2013), physical parameters (Wu et al., 2016), stability of a physical system (Groth et al., 2018), or involve question-answering for the agent to recognize conceptual differences (Weitnauer et al., 2023; Yi et al., 2019), or describe a physical scenario (Ates et al., 2020). For interactive tasks, AI is required to control some objects in the environment to complete certain tasks based on physical commonsense, e.g., solving classical mechanics puzzles (Bakhtin et al., 2019), flying a bird to reach a target position (Xue et al., 2023), and using a tool (Allen et al., 2020). Two works closely related to ours are Girdhar et al. (2020) and de Silva et al. (2020). Girdhar et al. (2020) introduce a forward prediction model to aid physical reasoning but do not address what is learned in the prediction model. de Silva et al. (2020) attempt to discover universal physical laws from data with abstracted internal states and human expertise introduced in the dynamic model design. Some works also aim to incorporate physical knowledge into video generation models through manually designed modules (Cao et al., 2024). In contrast, we focus on recovering physical laws from raw observation without any human priors, akin to a newborn baby.
Appendix B Latent Video Diffusion Model
B.1 Diffusion preliminaries
@@
where the target can be the corrupted noise , the original video , or the velocity between data and noise. Following (Salimans & Ho, 2022), we use the velocity prediction to train the video diffusion models, denoted by
Also, to eliminate the training and inference gap and ensure a zero signal-to-noise ratio at the final timestep, following (Lin et al., 2024), we set to when .
B.2 VAE Architecture and Pretrain
We commence with the structure of the SD1.5-VAE, retaining the majority of the original 2D convolution, group normalization, and attention mechanisms on the spatial dimensions. To inflate this structure into a spatial-temporal auto-encoder, we convert the final few 2D downsample blocks of the encoder and the initial few 2D upsample blocks of the decoder into 3D ones, and employ multiple extra 1D layers to enhance temporal modeling. For all the downsample blocks where temporal downsampling is required, we replace all the original 2D downsample layers with re-initialized causal 3D downsample layers by adding causal paddings to the head of the frame sequence (Yu et al., 2023b) and introduce an additional causal 1D convolution layer after the original ResNetBlock. As for the decoder part, all the 2D Nearest-Interpolation operations are substituted by a 2D convolution layer and a channel-to-space transformation, which are specifically initialized to behave precisely the same as Nearest-Interpolation operations before the first training step. For the discriminator part, we inherit the structure of the original 2D PatchGAN (Isola et al., 2016) discriminator used by SD1.5-VAE, and design a 3D PatchGAN discriminator based on the 2D version. Different from the generator module, we train the group of discriminators from scratch for the consideration of stability. Subsequently, all parameters of the (2+1)D-VAE are jointly trained with high-quality image and video data to preserve the capability of appearance modeling and to enable motion modeling. For the image dataset, we filter data samples from LAION-Aesthetics (Beaumont & Schuhmann, 2022), COYO (Byeon et al., 2022) and DataComp (Gadre et al., 2023) with high aesthetics and clarity to form a high-quality subset. As for the video dataset, we collect a high-quality subset from Vimeo-90K (Xue et al., 2017), Panda-70M (Wang et al., 2020) and HDVG (Wang et al., 2023). In the training process, we train the entire structure for 1M steps and only the random resized crop and random horizontal flip are applied in the data augmentation process.
B.3 VAE Reconstruction
@@
B.4 DiT Implementation Details
Following Sora (Brooks et al., 2024), we directly use the DiT architecture (Peebles & Xie, 2023) as the denoising model. The only modification is that we employ a 3D variant of RoPE to better fit the video data. Throughout the paper, we use a maximum learning rate of with cosine decay, a bach size of 256. All models are trained with the AdamW (Diederik & Ba, 2015; Loshchilov, 2017) optimizer, with , and the weight decay weight is equal to . Horizontal flipping is the only data augmentation if not specified.
Appendix C Detailed Experimental setup
C.1 Fundamental Physical Scenarios Data
@@
C.2 Combinatorial Experiments Evaluation Setup
To evaluate the generated videos, we performed human evaluations to determine the abnormal ratio, defined as the proportion of videos that violate physical laws as assessed by human intuition. We utilized 60 templates for training and reserved 10 unused templates for evaluation. For the in-template evaluation set, we sampled 2 videos from each of the 60 training templates, forming a total of 120 videos (excluded from training). For the out-of-template evaluation set, we sampled 10 videos from each of the 10 reserved templates, forming a total of 100 videos. We recruited 10 human evaluators to assess the videos. Each evaluator independently reviewed all videos and determined whether the physical processes depicted adhered to the physical laws. We average the abnormal rates over all the evaluators and report the result. Each evaluator received the following instruction: “For each video shown to you, all objects start with free fall. You should judge whether the physical processes in the video obey physical laws based on your intuition. Select ‘Yes’ if the process adheres to physical laws and ‘No’ otherwise. Please make your judgment based on your first impression”.
Appendix D Experiments on SOTA Video Generation Models
In this section, we explore how the findings in this paper apply to open-source pretrained models, such as CogVideo (Hong et al., 2022) and Stable Video Diffusion (SVD) (Blattmann et al., 2023).
D.1 Is the Prioritization Relevant to VAE?
To evaluate how much the prioritization order color > size > velocity > shape is influenced by the explicit instantiation of the video VAE used in our work, we perform experiments with alternative architectures. Specifically, we replace the VAE from our primary experiments with the VAE from the recently released CoGVideo and rerun the tests. By freezing the CoGVideo VAE and training only the diffusion component, we find that the prioritization order remains consistent.
As detailed in Section 5.3, we designed experiments to compare the prioritization of velocity and shape. Videos of balls with low velocities and squares with high velocities were used for training. During testing, we evaluated scenarios where a ball with high velocity transformed into a square immediately after the conditioned frames, and a square with low velocity transitioned into a ball. Across 1,400 test cases, no exceptions were observed. The results are identical to those presented in Figure 8 (3), confirming that the prioritization order velocity shape holds. For velocity and size, as shown in Figure 12 (1), the model consistently maintained the initial size and velocity for most test cases, even beyond the training distribution. However, a slight preference for size over velocity was observed, particularly when radius and velocity values were at extreme ends (top-left of the plot). These results confirm that size holds a higher priority than velocity. Finally, for the prioritization between color and velocity, we trained the model on high-speed blue balls and low-speed red balls. During testing, high-speed red balls appeared much slower than their conditioned velocity, while no balls in the test set changed their color. As illustrated in Figure 12 (2), this demonstrates that color is prioritized over velocity. In summary, our experiments with the CoGVideo VAE confirm the robustness of the prioritization order color > size > velocity > shape, demonstrating that these findings generalize across different VAE pretrain.
D.2 Can pretrained models learn physical laws?
The primary goal of this paper is to answer the fundamental question: “Can a video generation model learn physical laws?” Existing pretrained models are built on large-scale datasets sourced from publicly available internet content. However, the extent to which these datasets contain videos about physical laws or structured physical interactions remains unknown. This lack of transparency leads to critical concern - contamination of evaluation data: There is a risk that pretrained models may have been exposed to data similar to our evaluation experiments during training. This potential overlap undermines the validity of assessing their ability to learn physical reasoning. To address these concerns, we chose to train a model from scratch, which provides complete control over the training data. This approach allows us to precisely ensure that the results stem from the model’s actual learning process rather than potential pretraining data overlap or contamination. This methodology is supported by recent research into the mechanisms of language models, where training relatively small models from scratch has proven effective in isolating and analyzing reasoning processes more rigorously.
To assess whether the inability to learn physical laws is specific to our setup or extends to open-source pretrained models, we conducted experiments using the Stable Video Diffusion (SVD) model on a representative uniform motion task. We fine-tuned SVD for 300k steps, following the same training setup as DiT-B in our experiments. The results, summarized in Table 4, reveal several key observations. First, SVD’s VAE demonstrates a very small reconstruction error, indicating that its VAE component effectively preserves kinematic information. However, the diffusion component of SVD, despite pretraining, performs worse than DiT-B in both in-distribution (ID) and out-of-distribution (OOD) predictions. We attribute this to the considerable gap between the pretraining dataset (internet videos) and the simplified kinematic test scenarios used in this evaluation. Notably, the OOD error for SVD is an order of magnitude larger than its ID error, a trend consistent with DiT-B. This finding reinforces the robustness of our conclusion that pretraining alone does not address the inability of video generation models to generalize to OOD scenarios, a limitation that persists across architectures, including open-source pretrained models like SVD.
| Model | ID Error | OOD Error |
| GT | 0.0099 | 0.0104 |
| DiT-B (from scratch, ours) | 0.0138 | 0.3583 |
| SVD-VAE-Recon | 0.0103 | 0.0107 |
| SVD-Finetune | 0.0505 | 0.9081 |
Appendix E More Experiments and Discussions
E.1 Can Language and Numerics Aid in Learning Physical Laws?
@@
As discussed in Section 3, video generation based solely on image frames fails to learn physical laws, showing significant prediction errors in OOD scenarios, despite containing all the necessary information. In reinforcement learning, numerical values (e.g., states) are often used as conditions for world models (Hafner et al., 2023), and language representations have shown generalization capabilities in LLMs (Riveland & Pouget, 2024). This raises the question: can additional multimodal inputs, such as numerics and text, improve video prediction and capture physical laws? We experimented with collision scenarios and DiT-B models, adding two variants: one conditioned on vision and numerics, and the other on vision and text. For numeric conditioning, we map the state vectors to embeddings and add layer-wise features to video tokens. For text, we converted initial physical states into natural language descriptions, obtained text embeddings using a T5 encoder, and then added a cross-attention layer to aggregate textual representations for video tokens.
@@
E.2 Continuous Experiment for Pairwise Comparison
We also conducted attribute comparison experiments using continuous values for velocity and size to quantify these prioritizations. For velocity vs. size, the combinatorial generalization performance is surprisingly good. The model effectively maintains the initial size and velocity for most test cases beyond the training distribution. However, a slight preference for size over velocity is noted, particularly with extreme radius and velocity values (top left and bottom right in Figure 14 (1)). In Figure 14 (2), color can be combined with size most of the time. Conversely, for color vs. velocity in Figure 14 (3), high-speed blue balls and low-speed red balls are used for training. At test time, low-speed blue balls appear much faster than their conditioned velocity. No ball in the test set changes its color, indicating that color is prioritized over velocity. The above analysis is consistent with the conclusion drawn from the binary attribute comparisons: color > size > velocity > shape.
To confirm that the shift observed in Figure 14 is due in fact to attribute prioritization and not to general behavior, we performed control experiments corresponding to the main experiments shown in panels (2) and (3) of Figure 14. These experiments were designed without an attribute conflict (that is, only the red color was used). As shown in Figure 15, even in extreme cases, the predictions remained accurate without any change. This result indicates that the drift observed in Figure 14 arises specifically from the attribute conflict and the presence of attribute prioritization.
E.3 Principle Behind Data Retrieval in the Diffusion Model
@@
Since the diffusion model is trained by minimizing the loss associated with predicting VAE latent, we hypothesize that the prioritization may be related to the distance in VAE latent space (though we use pixel space here for a clearer illustration) between the test conditions and the training set. Intuitively, when comparing color and shape as in Figure 14 (1), a change in shape from a ball to a rectangle results in minor variation of the pixels, primarily at the corners. In contrast, a color change from blue to red causes a more significant pixel difference. Thus, the model tends to preserve color while allowing the shape to vary. From the perspective of pixel variation, the prioritization of color > size > velocity > shape can be explained by the extent of pixel change associated with each attribute. Changes in color typically result in large pixel variations because it affects nearly every pixel on its surface. In contrast, changes in size modify the number of pixels, but do not drastically alter the individual pixels’ values. Velocity affects pixel positions over time, leading to moderate variation as the object shifts, while shape changes often involve only localized pixel adjustments, such as at edges or corners. Therefore, the model prioritizes color because it causes the most significant pixel changes, while shape changes are less impactful in terms of pixel variation.
@@
To further validate this hypothesis, we designed a variant experiment comparing color and shape, as shown in Figure 16. In this case, we use a blue ball and a red ring. For the ring to transform into the ball without changing color, it would need to remove the ring’s external color, turning it into a blank space, and then fill the internal blank space with the ball’s color, resulting in significant pixel variation. Interestingly, in this scenario, unlike the previous experiments shown in Figure 14 (1), the prioritization of color ¿ shape does not hold. The red ring can transform into either a red ball or a blue ring, as demonstrated by the examples. This observation suggests that the model’s prioritization may indeed depend on the complexity of the pixel transformations required for each attribute change. Future work could explore more precise measurements of these variations in pixel or VAE latent space to better understand the model’s training data retrieval process.
E.4 Failure Cases in Combinatorial Generalization
@@
In Figure 11, we present successful cases of spatial and temporal combinatorial generalization, where the trained model can combine these two types of events in spatial and temporal dimensions to generate videos under unseen conditions. However, the model does not always succeed in performing such compositions, and here we illustrate some failure cases. As shown in Figure 17, when the training set lacks a red ball in a bounce event, model sometimes struggles. In the model-generated video, the red ball sometimes disappears after a collision. This likely stems from the model’s data retrieval mechanism: since the training set does not include a red ball in a collision scenario, when collision happens, the model retrieves similar training cases without the red ball, causing it to vanish post-collision. In summary, while combinatorial generalization allows the diffusion model to generate novel videos by composing spatial and temporal segments from the training set, its reliance on data retrieval limits its effectiveness. As a result, the model may produce unrealistic outcomes by retrieving and combining segments without understanding the underlying rules.
Appendix F Comparison with ID/OOD Generalization Works
ID/OOD generalization (Schott et al., 2021) in machine learning has been extensively discussed in foundational works. These discussions highlight the fundamental challenges of OOD generalization in machine learning. However, when applied to specific settings with unique data forms, distributions, and model inductive biases, it exhibits nuanced behaviors that remain scientifically significant. For instance, similar debates have emerged in the study of large language models (LLMs), where considerable attention has been devoted to determining whether models learn underlying rules or rely on case-based memorization for performing arithmetic (Hu et al., 2024). These insights have spurred advancements in rule-based reasoning approaches within the LLM community. Similarly, our work investigates how video generation models handle generalization within their own domain.
Our findings go beyond the simplistic observation that ID scenarios succeed while OOD scenarios fail. Instead, we provide deeper insights and discoveries specific to video generation:
1.
Addressing the Debate on Learning Physical Laws in Video Generation Models: Through systematic experiments, we provide a clear answer to whether physical laws can be learned by scaling video generation models. This has been a topic of debate in the video generation community. For instance, OpenAI’s Sora Technical Report (Brooks et al., 2024) suggests that video generation models can learn rules and act as world simulators, implying they might generalize universal laws. Our findings challenge this assumption, showing that current video generation models fail to learn physical laws as universal rules.
2.
Scaling Guidance for Combinatorial Generalization: We demonstrate the importance of combinatorial generalization and identify scaling laws for improving generalization in video generation models. Our findings highlight that increasing the diversity of combinations in the training data is more effective for achieving realistic physics than merely scaling the data volume or model size. This offers actionable guidance for building video generation models that better capture physical realism.
3.
Revealing the Generalization Mechanism and Understanding the Boundaries of Video Generation Models: We uncover how video generation models generalize, primarily relying on referencing similar training examples rather than learning underlying universal principles. This provides a deeper understanding of their limitations and biases in representing physical phenomena.
These findings transcend traditional ID/OOD analyses, offering a more detailed understanding of the limitations and potential of video generation models.
Appendix G Visualization